# Pricing Options on Forwards in Function-Valued Affine Stochastic Volatility Models

### v1

@@ Appendix B Numerical method for solving the Riccati equations in the option pricing of Wishart model
In the characteristic function and the option pricing equation, one has to compute the two quantities $\langle X_{t},q_{1}(T-t,u)\rangle,\operatorname{Tr}(Y_{t}q_{2}(T-t,u))$ and P(Tt,u)P(T-t,u). Specially, in implementation we define $t=0$ and $Y_{0}=Q$. Note that $q_{1}(t,u)$ has the same solution as $\psi_{1}(t,u)$ in equation (4.5), and hence $\langle X_{t},q_{1}(T-t,u)\rangle$ can be easily calculated. Moreover, the Riccati equation for $P(t,u)$ is given by
$\frac{\partial P}{\partial t}(t,u)=m\operatorname{Tr}(Qq_{2}(t,u)).$
Note that
$\operatorname{Tr}(Qq_{2}(t,u))$ $=\sum_{n=1}^{\infty}\langle Qq_{2}(t,u)f_{n},f_{n}\rangle$
$=\sum_{n=1}^{\infty}\langle\frac{1}{2}\sum_{k=1}^{\infty}\frac{1}{k^{2}}\langle q_{2}(t,u)f_{n},f_{k}\rangle f_{k},f_{n}\rangle$
$=\frac{1}{2}\sum_{n=1}^{\infty}\frac{1}{n^{2}}\langle q_{2}(t,u)f_{n},f_{n}\rangle.$
Therefore, in order to obtain the characteristic function and the option pricing equation, one has to compute $\langle q_{2}(t,u)f_{n},f_{n}\rangle,n=1,2,\ldots$. Consider the more general case $\langle q_{2}(t,u)f_{i},f_{j}\rangle,i,j=1,2,\ldots$. Recall the Riccati equations for $q_{2}(t,u)$ in the Wishart model are given by
$\frac{\partial q_{2}}{\partial t}(t,u)=Qq_{2}(t,u)+q_{2}(t,u)Q-\frac{1}{2}(D^{\frac{1}{2}}q_{1}(t,u))-\frac{1}{2}(q_{2}(t,u)+q_{2}^{T}(t,u))Q(q_{2}(t,u)+q_{2}^{T}(t,u))$
Discreterize this dynamic, one obtains
$q_{2}(t+\delta,u)$ $=q_{2}(t,u)+\delta(Qq_{2}(t,u)+q_{2}(t,u)Q)-\frac{\delta}{2}(D^{\frac{1}{2}}q_{1}(t,u))^{\otimes 2}$ (B.1)
$-\frac{\delta}{2}\left(q_{2}(t,u)+q_{2}^{T}(t,u))Q(q_{2}(t,u)+q_{2}^{T}(t,u))\right).$
According to equation (B.1), $\langle q_{2}(t,u)f_{i},f_{j}\rangle$ has the iteration
$\langle q_{2}(t+\delta,u)f_{i},f_{j}\rangle-\langle q_{2}(t,u)f_{i},f_{j}\rangle$ (B.2)
$=\delta\langle Qq_{2}(t,u)f_{i},f_{j}\rangle+\delta\langle q_{2}(t,u)Qf_{i},f_{j}\rangle$
$-\frac{\delta}{2}\langle(D^{\frac{1}{2}}q_{1}(t,u))^{\otimes 2}f_{i},f_{j}\rangle-\frac{\delta}{2}\langle(q_{2}(t,u)+q_{2}^{T}(t,u))Q(q_{2}(t,u)+q_{2}^{T}(t,u))f_{i},f_{j}\rangle.$
We will compute the four terms on right hand side of the equation (B.2). First of all, note that
$Qq_{2}(t,u)f_{j}=\frac{1}{2}\sum_{n=1}^{\infty}\frac{1}{n^{2}}\langle q_{2}(t,u)f_{i},f_{n}\rangle f_{n}.$
One obtains
$\langle Qq_{2}(t,u)f_{i},f_{j}\rangle=\frac{1}{2}\langle\sum_{n=1}^{\infty}\frac{1}{n^{2}}\langle q_{2}(t,u)f_{i},f_{n}\rangle f_{n},f_{j}\rangle=\frac{1}{2}\frac{1}{j^{2}}\langle q_{2}(t,u)f_{i},f_{j}\rangle.$ (B.3)
Similarly,
$\langle q_{2}(t,u)Qf_{i},f_{j}\rangle=\frac{1}{2}\frac{1}{i^{2}}\langle q_{2}(t,u)f_{i},f_{j}\rangle.$ (B.4)
For $\langle(D^{\frac{1}{2}}q_{1}(t,u))^{\otimes 2}f_{i},f_{j}\rangle$, note that
$(D^{\frac{1}{2}}q_{1}(t,u))^{\otimes 2}f_{i}$ $=\frac{1}{2}(\nu+\mathrm{i}\lambda)^{2}(\sum_{n=1}^{\infty}\frac{1}{n}f_{n}(t+\vartheta)\langle f_{n},f_{i}\rangle)(\sum_{n=1}^{\infty}\frac{1}{k}f_{k}(t+\vartheta)f_{k})$
$=\frac{1}{2}(\nu+\mathrm{i}\lambda)^{2}\frac{1}{i}f_{i}(t+\vartheta)(\sum_{n=1}^{\infty}\frac{1}{k}f_{k}(t+\vartheta)f_{k}).$
It follows that
$\langle(D^{\frac{1}{2}}q_{1}(t,u))^{\otimes 2}f_{i},f_{j}\rangle=\frac{1}{2}(\nu+\mathrm{i}\lambda)^{2}\frac{1}{ij}f_{i}(t+\vartheta)f_{j}(t+\vartheta).$ (B.5)
Finally, to compute $\langle(q_{2}(t,u)+q_{2}^{T}(t,u))Q(q_{2}(t,u)+q_{2}^{T}(t,u))f_{i},f_{j}\rangle$, we first compute the quantity $(q_{2}(t,u)+q_{2}^{T}(t,u))Q(q_{2}(t,u)+q_{2}^{T}(t,u))f_{i}$ as below:
$(q_{2}(t,u)+q_{2}^{T}(t,u))Q(q_{2}(t,u)+q_{2}^{T}(t,u))f_{i}$
$=\frac{1}{2}(q_{2}(t,u)+q_{2}^{T}(t,u))\sum_{k=1}^{\infty}\frac{1}{k^{2}}\langle q_{2}(t,u)f_{i}+q_{2}^{T}(t,u)f_{i},f_{k}\rangle f_{k}$
$=\frac{1}{2}(q_{2}(t,u)+q_{2}^{T}(t,u))\sum_{k=1}^{\infty}\frac{1}{k^{2}}\left(\langle q_{2}(t,u)f_{i},f_{k}\rangle+\langle q_{2}(t,u)f_{k},f_{i}\rangle\right)f_{k}$
$=\frac{1}{2}\sum_{k=1}^{\infty}\frac{1}{k^{2}}\left(\langle q_{2}(t,u)f_{i},f_{k}\rangle+\langle q_{2}(t,u)f_{k},f_{i}\rangle\right)(q_{2}(t,u)+q_{2}^{T}(t,u))f_{k}$
Consequently
$\langle(q_{2}(t,u)+q_{2}^{T}(t,u))Q(q_{2}(t,u)+q_{2}^{T}(t,u))f_{i},f_{j}\rangle$ (B.6)
$=\frac{1}{2}\sum_{k=1}^{\infty}\frac{1}{k^{2}}\left(\langle q_{2}(t,u)f_{i},f_{k}\rangle+\langle q_{2}(t,u)f_{k},f_{i}\rangle\right)\langle(q_{2}(t,u)+q_{2}^{T}(t,u))f_{k},f_{j}\rangle$
$=\frac{1}{2}\sum_{k=1}^{\infty}\frac{1}{k^{2}}\left(\langle q_{2}(t,u)f_{i},f_{k}\rangle+\langle q_{2}(t,u)f_{k},f_{i}\rangle\right)\left(\langle q_{2}(t,u)f_{k},f_{j}\rangle+\langle q_{2}(t,u)f_{j},f_{k}\rangle\right)$
$=\frac{1}{2}\sum_{k=1}^{\infty}\frac{1}{k^{2}}(\langle q_{2}(t,u)f_{i},f_{k}\rangle\langle q_{2}(t,u)f_{k},f_{j}\rangle+\langle q_{2}(t,u)f_{i},f_{k}\rangle\langle q_{2}(t,u)f_{j},f_{k}\rangle$
$\quad+\langle q_{2}(t,u)f_{k},f_{i}\rangle\langle q_{2}(t,u)f_{k},f_{j}\rangle+\langle q_{2}(t,u)f_{k},f_{i}\rangle\langle q_{2}(t,u)f_{j},f_{k}\rangle)$
Plugging the equations (B.3), (B.4), (B.5) and (B.6) into equation (B.2), one obtains
$\langle q_{2}(t+\delta,u)f_{i},f_{j}\rangle$ $=\langle q_{2}(t,u)f_{i},f_{j}\rangle+\frac{\delta}{2}\frac{1}{j^{2}}\langle q_{2}(t,u)f_{i},f_{j}\rangle+\frac{\delta}{2}\frac{1}{i^{2}}\langle q_{2}(t,u)f_{i},f_{j}\rangle$ (B.7)
$-\frac{\delta}{4}(\nu+\mathrm{i}\lambda)^{2}\frac{1}{ij}f_{i}(t+\vartheta)f_{j}(t+\vartheta)$
$-\frac{\delta}{4}\sum_{k=1}^{\infty}\frac{1}{k^{2}}(\langle q_{2}(t,u)f_{i},f_{k}\rangle\langle q_{2}(t,u)f_{k},f_{j}\rangle+\langle q_{2}(t,u)f_{i},f_{k}\rangle\langle q_{2}(t,u)f_{j},f_{k}\rangle$
$+\langle q_{2}(t,u)f_{k},f_{i}\rangle\langle q_{2}(t,u)f_{k},f_{j}\rangle+\langle q_{2}(t,u)f_{k},f_{i}\rangle\langle q_{2}(t,u)f_{j},f_{k}\rangle).$
In implementation, one has to truncate the infinite sum on the right hand side of equation (B.7) by the chosen dimensionality $n$. Therefore, the iteration in (B.7) is approximated by
$\langle q_{2}(t+\delta,u)f_{i},f_{j}\rangle$ $\approx\langle q_{2}(t,u)f_{i},f_{j}\rangle+\frac{\delta}{2}\frac{1}{j^{2}}\langle q_{2}(t,u)f_{i},f_{j}\rangle+\frac{\delta}{2}\frac{1}{i^{2}}\langle q_{2}(t,u)f_{i},f_{j}\rangle$ (B.8)
$-\frac{\delta}{4}(\nu+\mathrm{i}\lambda)^{2}\frac{1}{ij}f_{i}(t+\vartheta)f_{j}(t+\vartheta)$
$-\frac{\delta}{4}\sum_{k=1}^{n}\frac{1}{k^{2}}(\langle q_{2}(t,u)f_{i},f_{k}\rangle\langle q_{2}(t,u)f_{k},f_{j}\rangle+\langle q_{2}(t,u)f_{i},f_{k}\rangle\langle q_{2}(t,u)f_{j},f_{k}\rangle$
$+\langle q_{2}(t,u)f_{k},f_{i}\rangle\langle q_{2}(t,u)f_{k},f_{j}\rangle+\langle q_{2}(t,u)f_{k},f_{i}\rangle\langle q_{2}(t,u)f_{j},f_{k}\rangle).$
for $i,j=1,\ldots,n$. Denote
$F(t)=\begin{bmatrix}\langle q_{2}(t,u)f_{1},f_{1}\rangle&\langle q_{2}(t,u)f_{1},f_{2}\rangle&\cdots&\langle q_{2}(t,u)f_{1},f_{n}\rangle\\ \langle q_{2}(t,u)f_{2},f_{1}\rangle&\langle q_{2}(t,u)f_{2},f_{2}\rangle&\cdots&\langle q_{2}(t,u)f_{2},f_{n}\rangle\\ \vdots&\vdots&\vdots&\vdots\\ \langle q_{2}(t,u)f_{n},f_{1}\rangle&\langle q_{2}(t,u)f_{n},f_{2}\rangle&\cdots&\langle q_{2}(t,u)f_{n},f_{n}\rangle,\end{bmatrix}$
$B=[\frac{1}{1}f_{1}(t+\vartheta),\frac{1}{2}f_{2}(t+\vartheta),\ldots,\frac{1}{n}f_{n}(t+\vartheta)],$
and
$N=\begin{bmatrix}\frac{1}{1^{2}}&&&\\ &\frac{1}{2^{2}}&&\\ &&\ddots&\\ &&&\frac{1}{n^{2}}\end{bmatrix}$
The Equation (B.8) can be reformed to a matrix form
$F(t+\delta)$ $=F(t)+\frac{\delta}{2}F(t)N+\frac{\delta}{2}NF(t)-\frac{\delta}{4}(\nu+\mathrm{i}\lambda)^{2}B^{T}B$ (B.9)
$-\frac{\delta}{4}(F(t)NF(t)+F(t)NF(t)^{T}+F(t)^{T}NF(t)+F(t)^{T}NF(t)^{T})$
This allows us to iteratively compute $F(t)$ with initial values $F(0)=0$. Therefore, the desired value $\operatorname{Tr}(Qq_{2}(t,u))$ can be computed by obtaining $\langle q_{2}(t,u)f_{k},f_{k}\rangle,k=1,\ldots,n$ from the diagonal of $F(t)$.

### v2

@@ Appendix B Numerical method for solving the Riccati equations in the option pricing of Wishart model
In the characteristic function and the option pricing equation, one has to compute the quantities $\langle X_{0},q_{1}(T,u)\rangle_{w}$, $\operatorname{Tr}(Y_{0}q_{2}(T,u))$ and $P(T,u)$. The first term is explicit because $q_{1}(t,u)$ solves the same linear equation as $\psi_{1}(t,u)$ in (4.5). The scalar Riccati term satisfies
$\frac{\partial P}{\partial t}(t,u)=n\operatorname{Tr}(Qq_{2}(t,u)),\qquad P(0,u)=0.$
For the numerical Wishart experiment of Section 4, we work with the diagonal parameterization
$Q$ $=\sum_{k\geq 1}q_{k}\,f_{k}\otimes f_{k},\qquad\mathbb{A}=-\sum_{k\geq 1}a_{k}\,f_{k}\otimes f_{k},\qquad D=\frac{1}{2}\sum_{k\geq 1}\frac{1}{k^{2}}\,f_{k}\otimes f_{k},$
$\Upsilon^{\vartheta}$ $=-\tfrac{1}{2}S^{*}(\vartheta)h_{0},\qquad D^{1/2}=\frac{1}{\sqrt{2}}\sum_{k\geq 1}\frac{1}{k}\,f_{k}\otimes f_{k},$
with $q_{k}>0$ and $a_{k}>0$. For $u_{1}=(\nu+\mathrm{i}\lambda)u_{\vartheta}$, define
$b_{k}(t,u)$ $:=\langle D^{1/2}q_{1}(t,u),f_{k}\rangle_{w}=\frac{\nu+\mathrm{i}\lambda}{\sqrt{2}\,k}\,f_{k}(t+\vartheta),$
$c_{k}^{\vartheta}$ $:=\langle D^{1/2}\Upsilon^{\vartheta},f_{k}\rangle_{w}=-\frac{1}{2\sqrt{2}\,k}\,\langle f_{k},S^{*}(\vartheta)h_{0}\rangle_{w}.$
Moreover, let
$F_{ij}(t,u):=\langle q_{2}(t,u)f_{i},f_{j}\rangle_{w},\qquad i,j\geq 1.$
Note that the coefficients $F_{ij}$ need not satisfy $F_{ij}=F_{ji}$, because the mixed term in (2.12) is not symmetry preserving. Since $q_{2}$ solves the Wishart Riccati equation associated with (2.12),
$\partial_{t}q_{2}(t,u)$ $=q_{2}(t,u)\mathbb{A}^{*}+\mathbb{A}q_{2}(t,u)-\tfrac{1}{2}(D^{1/2}q_{1}(t,u))^{\otimes 2}-D^{1/2}q_{1}(t,u)\otimes D^{1/2}\Upsilon^{\vartheta}$
$\quad-\tfrac{1}{2}(q_{2}(t,u)+q_{2}(t,u)^{T})Q(q_{2}(t,u)+q_{2}(t,u)^{T}),$
we obtain, for all $i,j\geq 1$,
$\frac{\partial}{\partial t}F_{ij}(t,u)$ $=-(a_{i}+a_{j})\,F_{ij}(t,u)-\frac{1}{2}\,b_{i}(t,u)b_{j}(t,u)-b_{i}(t,u)c_{j}^{\vartheta}$ (B.1)
$\quad-\frac{1}{2}\sum_{k=1}^{\infty}q_{k}\,\big(F_{ik}(t,u)+F_{ki}(t,u)\big)\big(F_{kj}(t,u)+F_{jk}(t,u)\big).$
To implement the scheme, we truncate the system to $H_{n}=\mathrm{span}\{f_{1},\dots,f_{n}\}$. Define
$\mathbf{F}_{n}(t,u):=\big(F_{ij}(t,u)\big)_{1\leq i,j\leq n},\qquad\mathbf{A}_{n}:=\mathrm{diag}(a_{1},\dots,a_{n}),\qquad\mathbf{Q}_{n}:=\mathrm{diag}(q_{1},\dots,q_{n}),$
$\mathbf{b}_{n}(t,u):=\big(b_{1}(t,u),\dots,b_{n}(t,u)\big)^{\top},\qquad\mathbf{c}_{n}^{\vartheta}:=\big(c_{1}^{\vartheta},\dots,c_{n}^{\vartheta}\big)^{\top}.$
An explicit Euler step with mesh size $\delta>0$ yields the recursion
$\mathbf{F}_{n}(t+\delta,u)$ $\approx\mathbf{F}_{n}(t,u)-\delta\big(\mathbf{A}_{n}\mathbf{F}_{n}(t,u)+\mathbf{F}_{n}(t,u)\mathbf{A}_{n}\big)$ (B.2)
$\quad-\frac{\delta}{2}\,\mathbf{b}_{n}(t,u)\mathbf{b}_{n}(t,u)^{\top}-\delta\,\mathbf{b}_{n}(t,u)(\mathbf{c}_{n}^{\vartheta})^{\top}$
$\quad-\frac{\delta}{2}\big(\mathbf{F}_{n}(t,u)+\mathbf{F}_{n}(t,u)^{\top}\big)\mathbf{Q}_{n}\big(\mathbf{F}_{n}(t,u)+\mathbf{F}_{n}(t,u)^{\top}\big).$
The initial condition is
$\mathbf{F}_{n}(0,u)=\big(\langle u_{2}f_{i},f_{j}\rangle_{w}\big)_{1\leq i,j\leq n}.$
In the first-moment experiment of Section 4, we take $u_{2}=0$, hence $\mathbf{F}_{n}(0,u)=0$.
The trace term is then approximated by
$\operatorname{Tr}(Qq_{2}(t,u))\approx\sum_{k=1}^{n}q_{k}\,F_{kk}(t,u)=\operatorname{Tr}(\mathbf{Q}_{n}\mathbf{F}_{n}(t,u)),$
and therefore
$P_{n}(t+\delta,u)\approx P_{n}(t,u)+\delta\,n\,\operatorname{Tr}(\mathbf{Q}_{n}\mathbf{F}_{n}(t,u)),\qquad P_{n}(0,u)=0.$
Together with the explicit formula for $q_{1}$, this gives the finite-rank transform approximation used in the Wishart numerical experiment. In Section 4 we further specialize to $q_{k}=a_{k}=1/k^{2}$.